Find the measure of an angle that is 26 degrees more than three times its complement
Answer by waynest(281)

angle 1 = x
angle 2= 3x + 26
x + (3x + 26) = 90
4x + 26 = 90
4x + 26 - 26 = 90 - 26
4x = 64
4x/4 = 64/4
x = 16
check:
16 + (3(16) + 26) = 90
16 + (48 + 26) = 90
16 + 74 = 90
90 = 90